# GDPR Elements, and Future Directions

### v3

@@ III-E Privacy Preservation
Bitcoin provides only weak identity privacy. To overcome this limitation, several privacy coins have been introduced, including Dash, Zerocoin, Monero, Zcash, Zether, Grin, Verge, and MimbleWimble, among others [60]. These cryptocurrencies aim to enhance transaction anonymity and protect user identities, making them relevant to privacy-focused, permissionless blockchains designed for identity privacy.
@@
Zerocoin (formerly Zcoin, XZC) is a proposal to extend Bitcoin with anonymous transactions. This extension would have functioned as a money laundering pool, where bitcoins were temporarily pooled in exchange for a temporary currency called zerocoins. Although laundering pools are already used by various currency laundering services, Zerocoin aimed to integrate this functionality at the protocol level, eliminating the need for trusted third parties. Ensured anonymity using cryptographic principles while still recording transactions within the Bitcoin blockchain [62].
@@
These privacy coins demonstrate how permissionless blockchains can be leveraged for identity privacy. By integrating cryptographic techniques such as ring signatures, zero-knowledge proofs, and coin-mixing mechanisms, these blockchain solutions provide individuals with enhanced control over their biometric data, mitigating the risks associated with traceable transactions.
More sophisticated general privacy-preservation mechanisms such as differential privacy, to our knowledge, have not yet been adapted and studied to blockchain.
@@ III-F Literature Categorization
This comparison aims to identify the strengths and limitations of different approaches, providing insight into how blockchain enhances biometrics for the given integration purpose. By analyzing these technical attributes, researchers and practitioners can better understand the trade-offs involved in various implementations and identify potential areas for future improvement.
@@
[87] On-chain and off-chain Fingerprint, face Biometric authentication Permissionless Transparency, convenience N/S Decision tree
@@
[36] On-chain Finger vein Library management N/S Availability N/S N/S
@@ IV Blockchain and Biometrics: GDPR Elements
As mentioned above, in contrast to traditional centralized database systems, blockchains allow "selective data exchange" with certain stakeholders, which can reduce the chance of identity theft and privacy problems [92, 93]. However, it brings about many other problems regarding the applicability of the GDPR. The analysis below is categorized according to five relevant parameters: (i) biometric data storage, (ii) accountability, (iii) immutability, (iv) anonymization, and (v) data protection by design and default. These are further addressed in the following, taking into account the technical analysis.
@@ IV-A Biometric Data Storage
This parameter is related to the questions as to how and where the template is stored. Interdisciplinary research [94] in 2009 concluded that biometrics could be invasive to privacy or improve to privacy depending on the design and safeguards. In addition to the research analyzed above, blockchain surveys show that depending on the types of network and storage, these systems might have different levels of GDPR compliance [95]. These two storage criteria are often suggested as key determinants of GDPR compliance level, whereby a helpful distinction can be made based on on-chain and off-chain storage.
@@ IV-A1 Off-chain storage
Central storage solutions are not considered privacy-friendly, as observed by the European Data Protection Supervisor [96]. Off-chain storage can be decentralized and such strict decentralization methods are considered more privacy friendly [97] as the data controller cannot access the biometric templates and only receives the comparison result. However, data security is a crucial part of the data protection law, and this type of decentralization may also be susceptible to attacks. As explained in various reports, such as by the BSI, local storage of biometric data enabled by software, for example, rich iOS of a mobile handset, cannot be regarded secure [98]. ENISA further observed that multiple biometric templates can be stored in Touch or FaceID and can belong to more than one person [99].
A common misunderstanding in the technical papers analyzed is that off-chain biometric data storage is, by default, GDPR friendly. However, as explained, off-chain storage itself does not render the processing compliant or secure. This implies that even if the biometric data are stored off-chain, personal data linked to it, such as a user ID, public keys, and the content of the transactions, are usually still stored on-chain. In such cases, even if the public keys, e.g., hashed identifiers, change in every transaction, they still qualify as personal data under the GDPR, as the natural persons remain identifiable.
@@ IV-A2 On-chain storage
On-chain processing, that is, the processing "located, performed, run inside a blockchain system" [100] of the special categories as such, while risky for the reasons we will discuss in detail later, is not per se against the GDPR, as long as the processing is lawful. For instance, in cases where the explicit consent of the data subject or another legal and legitimate ground under Article 9 is relied on, and the other GDPR requirements are met, this processing by itself would not infringe the GDPR. However, on-chain processing of personal data is not recommended for the reasons explained in further detail.
@@ IV-B Accountability
As explained previously, one of the most critical problems identified in the legal literature is how to allocate responsibility for data processing activities in the face of the polycentric nature of blockchains (i.e., accountability) [101, 102, 1]. Although private blockchains are less inclined to have such problems, the real benefits of blockchains, namely transparency, can be seen in public blockchains. That is why it is not surprising that identity management systems based on blockchain (and biometrics) are positioned to run their projects on public blockchains. This can be seen as a general conceptual issue with respect to public blockchains and requires more legal guidance and research. Hence, unless more legal clarity is established on the issue, processing via public blockchains is risky, as the entity responsible for GDPR compliance is unclear.
@@ IV-C Immutability
Immutability as the main attribute of blockchain may compromise the privacy of data subjects, particularly where their biometric data are processed [107]. Article 17 GDPR gives effect to the demands of individuals (data subjects) to have their data deleted without undue delay. As we just stated, biometric traits are unchangeable and biometric templates are, in theory, irrevocable. Therefore, any blockchain design that envisions raw or hashed on-chain storage of biometric templates would exacerbate the data protection risks. However, the same article also defines the exceptions to the right that are important to keep in mind when experimenting with blockchains and personal data. These exceptions demonstrate that the right to be forgotten is not an absolute right, nor is the right to privacy and the right to the protection of personal data. These exceptions are as follows.
@@ IV-D Anonymity
Under GDPR, encrypted personal data are considered pseudonymous in principle, not anonymous. Pseudonymity refers to the cases where personal data are linked to unique identifiers, the processing of which is still subject to the GDPR. Article 29 Working Party states that ‘removing directly identifying elements in itself is not enough to ensure that identification of the data subject is no longer possible. It will often be necessary to take additional measures to prevent identification, again depending on the context and purposes of the processing for which the anonymized data are intended’ [108, 109]. However, most blockchain research analyzed uses the term anonymous, for the data that are hashed or encrypted. This approach might create an incorrect understanding and implementation of the GDPR, where the responsible entities consider that the processing operation is not within the scope of the GDPR. It is specifically risky when it comes to hashed biometric templates, since these data are considered personal data [110].
@@ IV-E Data Protection by Design and by Default
For all the four problems explained above, another very relevant principle is data protection by design and default provided by Article 25 (1) GDPR. This principle is highly inclusive and starts to apply when the data controller determines the means of processing. Arguably, the developers of a blockchain-based biometric system can also be responsible. However, there are many arguments against this idea [111, 112]. However, the principle requires the data controller to take into account the state of the art, the cost of implementation, the nature, scope, purposes of personal data processing, and the risks and severity for the rights and freedoms of the data subjects. Biometric data processing is by nature considered risky. However, as mentioned, blockchain is still emerging and thus is considered a new technology whose impact on fundamental rights is unknown. Therefore, the processing of biometric data via blockchain is considered an innovative use of a large-scale operation that involves sensitive data where several datasets are matched or combined, and shall be subject to a Data Protection Impact Assessment (DPIA), under Article 35 GDPR[109].
@@ V-A Blockchain and Biometrics: Future Directions
V-A1 Biometric data should not be stored on a public ledger
As discussed in section IV, before having clarified the legal uncertainty on accountability, it is currently advised not to store biometric data on public blockchains. A public blockchain is an immutable storage that is accessible to anyone [100]. This means that once the data are published on the blockchain, it remains there forever. Considering the fact that biometric data contain information about our permanent identity and they are irreplaceable in case they leak. There are no guarantees that encryption mechanisms or biometric template protection methods remain completely safe in the future. Biometric data are discouraged in any form on a public blockchain.
In addition, from an efficiency point of view, blockchain storage is expensive both computationally and economically. In particular, reference [17] concluded practically that storing biometric templates on-chain or as hashed data is not cost-efficient. In addition to that, references [30] and [32] resorted to IPFS to solve the efficiency problem. Although using a distributed filesystem offers availability, it does not provide the security benefits that a blockchain provides, such as trust. Approaches (ii) and (iii) in Table I are proposed to address this concern.
V-A2 The integration should not be used for real-time scenarios
@@ V-A4 The integration should have economic justification
The life cycle of permissionless blockchains depends on the service of the miners who maintain the system. It should be economically interesting for these nodes to carry out their maintenance. To this end, the network pays a reward as a fee to miners who allocate their computing resources to maintain the network. For example, in Ethereum, this fee is called gas, and the required amount of gas to run the instructions inside the network differs according to their computational complexity [26]. Therefore, running smart contracts on future frameworks that combine blockchain and biometrics should be economically justifiable.
@@ V-B Combining Blockchain and Biometrics: Use Cases
So far, the most attractive applications for integrating blockchain and biometrics have been biometric verification and identification. Although these are conventional and widespread biometric applications, they are not necessarily the most fruitful use cases. After delving into existing studies, we conclude that using blockchain was not strictly necessary for many of these applications and was just added because of the recent blockchain hype. For example, it is mentioned in reference [81] that it is not clear to them why the blockchain is better than any regular storage system, yet they still use it. In the following, we propose some highly potential use cases for integrating blockchain and biometrics considering the pros and cons discussed in the rest of the paper.
@@ V-B1 Using blockchain as computing power for biometrics
The first consensus protocol employed in the blockchain network is Proof-of-Work (PoW), used in the Bitcoin network [114]. PoW is a complicated and resource-intensive strategy to select the miner who wins the competition, being entitled to publish the next block on the network. The difficulty level of this competition is continuously adjusted by the network to maintain trust and preserve the immutability of the blockchain. Despite these advantages, PoW is not economically efficient or environmentally friendly as it consumes substantial electrical power. When a block is published, only one node (miner) is rewarded for allocating its computational power, while the resources of other nodes who lose the competition are wasted. To remedy this issue, Proof-of-Stake (PoS) has been proposed [115]. However, PoS also has its problems regarding security threats and stakes issues [43]. A fruitful use case that benefits both blockchain and biometrics concerning PoW might be to recruit or rent the computational power of miners. Biometric recognition can be a computationally expensive task, especially in relation to the model training phase. It is highly profitable to harness the computational power of miners in PoW-based blockchains to carry the burden of biometric computation for tasks that existed in different stages of a biometric system, such as training a biometric model, generating a protected template, and comparing biometric templates. In particular, it seems that the blockchain could be a profitable platform for both clients and model owners to operate a federated learning system [116] for biometrics. Therefore, in the proposed consensus algorithm, those nodes who dedicate their computational power to the benefits of the blockchain service for the biometric system are rewarded according to their contribution.
@@ V-B3 Using blockchain as a foundation for biometrics-based PKI mechanisms
Integrating biometrics with PKI is an emerging approach aimed at enhancing security and user authentication. However, the adoption of biometric-based PKI systems is not yet widespread. Traditional PKI relies on passwords or hardware tokens for authentication and is managed by centralized authorities, making it vulnerable to single points of failure and inefficiencies. Incorporating biometrics can improve security by linking cryptographic keys to unique physiological traits. In contrast, blockchain-based PKI decentralizes key management in a network of nodes, reducing dependence on a central authority and improving resilience[118]. Despite this potential, several factors contribute to its limited adoption.
@@
World Network (formerly Worldcoin) and Humanity Protocol are examples of biometric-based PKI systems that use blockchain technology. World Network employs the biometric iris scan depicted in Fig. 5 to verify human identity, creating private crypto tokens [120, 121] for applications such as universal basic income and fraud prevention in virtual communications[122]. (Note that iris biometrics can also be captured with standard smartphone cameras [123], and iris biometrics, like other biometrics, are also susceptible to potential attacks and vulnerabilities [124].) Similarly, the Humanity Protocol uses palm scans [125] to authenticate online accounts, addressing issues such as bots [126, 127, 128], fake accounts, and online fraud, using blockchain and biometrics for secure identity verification[129].

### v4

@@ III-E Privacy Preservation
Bitcoin provides only weak identity privacy. To overcome this limitation, several privacy coins have been introduced, including Dash, Zerocoin, Monero, Zcash, Zether, Grin, Verge, and MimbleWimble, among others [61]. These cryptocurrencies aim to enhance transaction anonymity and protect user identities, making them relevant to privacy-focused, permissionless blockchains designed for identity privacy, which is crucial for biometric systems.
@@
Zerocoin (formerly Zcoin, XZC) is a proposal to extend Bitcoin with anonymous transactions. This extension would have functioned as a money laundering pool, where bitcoins were temporarily pooled in exchange for a temporary currency called zerocoins. Although laundering pools are already used by various currency laundering services, Zerocoin aimed to integrate this functionality at the protocol level, eliminating the need for trusted third parties. This Ensures anonymity using cryptographic principles while still recording transactions within the Bitcoin blockchain [63].
@@
These privacy coins demonstrate how permissionless blockchains can be leveraged for identity privacy. By integrating cryptographic techniques such as ring signatures [66], zero-knowledge proofs [67], and coin-mixing mechanisms [68], these blockchain solutions provide individuals with enhanced control over their biometric data, mitigating the risks associated with traceable transactions. More sophisticated general privacy-preservation mechanisms such as differential privacy, to our knowledge, have not yet been adapted and studied to blockchain.
@@ III-F Literature Categorization
This comparison aims to identify the strengths and limitations of different approaches, providing insight into how blockchain enhances biometrics for the given integration purpose. By analyzing these technical attributes, researchers and practitioners can better understand the trade-offs involved in various implementations and identify potential areas for future improvement. Given that, only a few works include prototypes and even those lack standardized metrics (latency, throughput, storage cost), a direct quantitative comparison is not feasible. Accordingly, Table III presents a semi-quantitative aggregation, derived from the attributes cataloged in Table II.
@@
[36] On-chain Finger vein Library management N/S Availability N/S N/S
[91] Off-chain and On-chain N/S biometric authentication Permissionless Decentralized authentication, IdM without re-enrollment Ethereum Fuzzy Commitment Scheme (FCS)
[92] Off-chain and On-chain Fingerprint Authentication Permissioned Decentralized authentication Hyperledger Fabric Fuzzy Extractor, Fabric chain
@@
[93] Off-chain and on-chain Fingerprint, face Biometric authentication Permissionless Transparency, convenience N/S Decision tree
@@ IV Blockchain and Biometrics: GDPR Elements
Unlike centralized databases, blockchains enable selective data sharing that can mitigate identity theft and privacy risks [98, 99]. However, they raise several GDPR-related challenges, analyzed below across five key parameters: biometric data storage, accountability, immutability, anonymization, and data protection by design and default.
@@ IV-A Biometric Data Storage
This parameter is related to the questions as to how and where the template is stored. Interdisciplinary research [100] in 2009 concluded that biometrics could improve privacy depending on the design and safeguards. In addition to the research analyzed above, blockchain surveys show that depending on the types of network and storage, these systems might have different levels of GDPR compatibility with the right to data protection [101]. These two storage criteria are often suggested as key determinants of GDPR compatibility level, whereby a helpful distinction can be made based on on-chain and off-chain storage.
@@ IV-A1 Off-chain storage
Typically, central storage solutions are not considered privacy-friendly, as observed by the European Data Protection Supervisor [102]. Off-chain storage of biometric references can be decentralized and such strict decentralization methods are considered more privacy friendly [103] as the data controller cannot access the biometric templates and only receives the comparison result. However, data security is a crucial part of the data protection law, and this type of decentralization may also be susceptible to attacks. As explained in various reports, such as by the BSI, local storage of biometric data enabled by software, for example, rich OS of a mobile handset, cannot be regarded secure [104]. ENISA further observed that multiple biometric templates can be stored in Touch or FaceID and can belong to more than one person [105].
A common misunderstanding in the technical papers analyzed is that off-chain biometric data storage is, by default, GDPR-compliant. However, as explained, off-chain storage itself does not make the processing compliant or secure. This implies that even if the biometric data are stored off-chain, personal data linked to it, such as a user ID, public keys, and the content of the transactions, are usually still stored on-chain. In such cases, even if the public keys, e.g., hashed identifiers, change in every transaction, they still qualify as personal data under the GDPR, as the natural persons remain identifiable. Nevertheless, off-chain storage of personal data is generally recommended, a practice also supported by the recent guidelines from the EDPB [106].
@@ IV-A2 On-chain storage
On-chain processing, that is, the processing "located, performed, run inside a blockchain system" [107] of the special categories as such, while risky for the reasons we will discuss in detail later, is not per se against the GDPR, as long as the processing is lawful. For instance, in cases where the explicit consent of the data subject or another legal and legitimate ground under Article 9 is relied on, and the other GDPR requirements are met, this processing by itself would not infringe the GDPR.
@@ IV-B Accountability
As explained previously, one of the most critical problems identified in the legal literature is how to allocate responsibility for data processing activities in the face of the polycentric nature of blockchains (i.e., accountability) [108, 109, 1]. Although private blockchains are less inclined to have such problems, the real benefits of blockchains, namely transparency, can be seen in public blockchains. That is why it is not surprising that identity management systems based on blockchain (and biometrics) are positioned to run their projects on public blockchains. This can be seen as a general conceptual issue with respect to public blockchains and requires more legal guidance and research. The EDPB Guidelines acknowledge that participants’ roles vary based on the governance model. In permissionless blockchains, nodes can influence processing and key decisions, such as forking. When nodes act independently, the EDPB recommends forming a consortium or legal entity among nodes. This legal structure would serve as the controller for processing [106]. Hence, unless more legal clarity is established on the issue, processing via public blockchains is risky, as the entity responsible for GDPR compliance is unclear.
@@
TABLE III: Semi-quantitative summary of existing blockchain–biometric integration approaches.
Approach Type # Papers Common Blockchain Type Implementation Status GDPR Compliance Risk Scalability
Full On-chain Storage 2424 Bitcoin, Ethereum Mostly conceptual; few demos High risk Poor
Off-chain + Blockchain Anchor 1010 Ethereum (+ FCS), Hyperledger Fabric Multiple prototypes (Fabric; ETH dApp) Medium risk Moderate–Good
Hybrid (mixed) 33 Sidechains, Hyperledger Prototype level Medium risk Moderate
SSI / Interoperable Identity 44 Hyperledger, Ethereum Emerging pilots Low risk Good
Audit Trails / Logging 22 Mixed / not specified Conceptual Medium risk Good
Privacy-Enhancing Approaches 11 Mixed / not specified Conceptual only Low risk Poor
Notes: “FCS” stands for the Fuzzy Commitment Scheme, a cryptographic method for securely binding biometrics to random keys. “Blockchain Anchor” denotes architectures where only cryptographic references (e.g., hashes or commitments) are stored on-chain, while biometric data remain off-chain.
@@ IV-C Immutability
Immutability as the main attribute of blockchain may compromise the privacy of data subjects, particularly where their biometric data are processed [114]. Article 17 GDPR gives effect to the demands of individuals (data subjects) to have their data deleted without undue delay. As we just stated, biometric traits are unchangeable and biometric templates are, in theory, irrevocable. Therefore, any blockchain design that envisions raw or hashed on-chain storage of biometric templates would exacerbate the data protection risks. However, the same article also defines the exceptions to the right that are important to keep in mind when experimenting with blockchains and personal data. These exceptions demonstrate that the right to be forgotten is not an absolute right, nor is the right to privacy and the right to the protection of personal data. These exceptions are as follows:
@@ IV-D Anonymity
Under GDPR, encrypted personal data are considered pseudonymous in principle, not anonymous. Pseudonymity refers to the cases where personal data are linked to unique identifiers, the processing of which is still subject to the GDPR. Article 29 Working Party states that ’removing directly identifying elements in itself is not enough to ensure that identification of the data subject is no longer possible’. It will often be necessary to take additional measures to prevent identification, again depending on the context and purposes of the processing for which the anonymized data are intended’ [115, 116]. However, most of the blockchain research analyzed uses the term anonymous to refer to data that are merely hashed or encrypted. This approach might create an incorrect understanding and implementation of the GDPR, where the responsible entities consider that the processing operation is not within the scope of the GDPR. It is specifically risky when it comes to hashed biometric templates, since these data are considered personal data [117].
@@ IV-E Data Protection by Design and by Default
For all the four problems explained above, another very relevant principle is data protection by design and default provided by Article 25 (1) GDPR. This principle is highly inclusive and starts to apply when the data controller determines the means of processing. Arguably, the developers of a blockchain-based biometric system can also be responsible. However, there are many arguments against this idea [118, 119]. The principle nevertheless requires the data controller to take into account the state of the art, the cost of implementation, the nature, scope, purposes of personal data processing, and the risks and severity for the rights and freedoms of the data subjects. Biometric data processing is by nature considered risky. However, as mentioned, blockchain is still emerging and thus is considered a new technology whose impact on fundamental rights is unknown. Therefore, the processing of biometric data via blockchain is considered an innovative use of a large-scale operation that involves sensitive data where several datasets are matched or combined, and shall be subject to a Data Protection Impact Assessment (DPIA), under Article 35 GDPR[116].
@@ V-A Blockchain and Biometrics: Future Directions
V-A1 Complete Biometric data should not be stored on a public ledger
As discussed in section IV, before having clarified the legal uncertainty on accountability, it is currently advised not to store the complete biometric data on public blockchains. A public blockchain is an immutable storage that is accessible to anyone [107]. This means that once the data are published on the blockchain, it remains there forever. Considering the fact that biometric data contain information about our permanent identity and they are irreplaceable in case they leak. There are no guarantees that encryption mechanisms or biometric template protection methods remain completely safe in the future.
From an efficiency standpoint, blockchain storage is currently computationally and economically expensive. In particular, reference [17] concluded practically that storing biometric templates on-chain or as hashed data is not cost-efficient. In addition to that, references [30] and [32] resorted to IPFS to solve the efficiency problem. Although using a distributed filesystem offers availability, it does not provide the security benefits that a blockchain provides, such as trust.
Currently, approaches (ii) and (iii) in Table I are proposed to mitigate these concerns by keeping biometric data off-chain while anchoring proofs on-chain. However, a potential line of future research is to investigate GDPR-compliant designs that allow limited or partial on-chain storage of biometric data.
V-A2 Current integration is not suitable for real-time scenarios
@@
A promising research direction is to experiment with new-generation blockchain networks that overcome these latency and throughput limitations, such as Solana, which offer higher transaction speeds and lower confirmation times suitable for near real-time biometric integration.
@@ V-A4 The integration should have economic justification
The lifecycle of a permissionless blockchain depends on the miners who maintain the network. It should be economically interesting for these nodes to carry out their maintenance. To this end, the network pays a reward as a fee to miners who allocate their computing resources to maintain the network. For example, in Ethereum, this fee, known as gas, varies depending on the computational complexity of the operations executed within the network [26]. Therefore, running smart contracts on future frameworks that combine blockchain and biometrics should be economically justifiable.
@@ V-B Combining Blockchain and Biometrics: Use Cases
So far, the most attractive applications for integrating blockchain and biometrics have been biometric verification and identification. Although these are conventional and well-established biometric tasks, they are not necessarily the most promising use cases for this integration. After delving into the existing literature, we observed that in many cases, blockchain was not a fundamental requirement but rather incorporated due to its recent popularity. For instance, reference [85] itself notes uncertainty about why blockchain would be preferable to a conventional storage system when integrated with biometrics, yet still employs it. In the following, we propose several highly potential use cases where blockchain integration with biometrics could offer genuine technical or operational advantages, taking into account the strengths and limitations discussed throughout this paper.
@@ V-B1 Using blockchain as computing power for biometrics
The first consensus protocol employed in the blockchain network is Proof-of-Work (PoW), used in the Bitcoin network [121]. PoW is a complicated and resource-intensive strategy to select the miner who wins the competition, being entitled to publish the next block on the network. The difficulty level of this competition is continuously adjusted by the network to maintain trust and preserve the immutability of the blockchain. Despite these advantages, PoW is not economically efficient or environmentally friendly as it consumes substantial electrical power. When a block is published, only one node (miner) is rewarded for allocating its computational power, while the resources of other nodes who lose the competition are wasted. To remedy this issue, Proof-of-Stake (PoS) has been proposed [122]. However, PoS also has its problems regarding security threats and stakes issues [43]. A potential use case that benefits both blockchain and biometrics concerning PoW might be to leverage or rent out the computational power of miners. Biometric recognition can be a computationally expensive task, especially in relation to the model training phase. It is highly profitable to harness the computational power of miners in PoW-based blockchains to carry the burden of biometric computation for tasks that existed in different stages of a biometric system, such as training a biometric model, generating a protected template, and comparing biometric templates. In particular, it seems that the blockchain could be a profitable platform for both clients and model owners to operate a federated learning system [123] for biometrics. Therefore, in the proposed consensus algorithm, those nodes who dedicate their computational power to the benefits of the blockchain service for the biometric system are rewarded according to their contribution.
@@ V-B3 Using blockchain as a foundation for biometrics-based PKI mechanisms
Integrating biometrics with PKI is an emerging approach aimed at enhancing security and user authentication. However, the adoption of biometric-based PKI systems is not yet widespread. Traditional PKI relies on passwords or hardware tokens for authentication and is managed by centralized authorities, making it vulnerable to single points of failure and inefficiencies. Incorporating biometrics can improve security by linking cryptographic keys to unique physiological traits. In contrast, blockchain-based PKI decentralizes key management in a network of nodes, reducing dependence on a central authority and improving resilience[125]. Unlike existing biometric-PKI systems that rely on a CA often alongside Registration and Validation Authorities (RA/VA) to issue and manage certificates, blockchain-anchored biometric PKI decentralizes trust and certificate-status publication. In conventional designs (e.g., BioPKI [126]), biometrics derive or protect key material (e.g., deriving a biometric public key and reconstructing the private key from the user’s biometric), while enrollment, issuance, and verification remain CA/RA/VA-controlled, creating single points of failure and limited transparency. A blockchain-anchored architecture reduces reliance on any single CA by recording integrity-protected bindings (hashes/commitments) between identifiers, keys, and biometric-derived artifacts on a distributed ledger, enabling tamper-evident audit trails, transparent revocation, and interoperable identity proofs through smart contracts or DIDs.
Despite this potential, several factors contribute to its limited adoption.
@@
World Network (formerly Worldcoin) and Humanity Protocol are other examples of biometric-based PKI systems that use blockchain technology. World Network employs the biometric iris scan depicted in Fig. 6 to verify human identity, creating private crypto tokens [128, 129] for applications such as universal basic income and fraud prevention in virtual communications[130]. (Note that iris biometrics can also be captured with standard smartphone cameras [131], and iris biometrics, like other biometrics, are also susceptible to potential attacks and vulnerabilities [132].) Similarly, the Humanity Protocol uses palm scans [133] to authenticate online accounts, addressing issues such as bots [134, 135, 136], fake accounts, and online fraud, using blockchain and biometrics for secure identity verification[137].